OpenStudy (anonymous):

please T___T ......... help me find the answer to this complex number (x+iy)/(2x+3iy) and (5x-3iy)/(7x-2iy)

zepdrix (zepdrix):

\[\Large\rm \frac{x+\mathcal i y}{2x+3\mathcal i y}\]So uhhh, we want to multiply the top and bottom by the `complex conjugate` of the bottom.

zepdrix (zepdrix):

\[\Large\rm \frac{x+\mathcal i y}{2x+3\mathcal i y}\color{royalblue}{\left(\frac{2x-3\mathcal i y}{2x-3\mathcal i y}\right)}\]k? Understand how to multiply that out?

OpenStudy (anonymous):

oh yes i get it ... i thought it should be change in polar form then that will be the time to solve... thank you